# EVM-QuestBench: An Execution-Grounded Benchmark for Natural-Language Transaction Code Generation

### v1

@@ Appendix E Additional Analysis Figures
E.1 Composite Pattern Difficulty
@@
E.2 Step Overhead Distribution
@@
E.3 Atomic Subcategory Difficulty
@@ F.1 Quadrant Summary
Table 8 summarizes the four capability profiles based on AtomicAvg and CompositeAvg medians (46.5 and 73.0 respectively).
@@
High High 5 strong precision and workflows
High Low 3 strong precision, weaker workflows
Low High 4 weaker precision, stronger workflows
Low Low 5 weaker on both splits
Atomic only 3 composite split not executed
@@ F.2 Model Coverage and Quadrant Classification
Table 10 reports coverage, normalized averages, total score, and the quadrant label based on AtomicAvg and CompositeAvg medians.
@@
gemini-2.5-flash -0.40 24 1 3715.7
gemini-3-pro-preview -0.69 33 1 3756.8
gpt-5 -0.98 22 1 2443.3
deepseek-v3.2 +0.56 42 1 4088.3
gpt-oss-120b +0.44 38 1 3682.3
kimi-k2-thinking +0.71 49 2 3504.9
grok-4-fast +1.00 53 3 3357.5
claude-haiku-4.5 +2.49 89 4 2214.7
glm-4.6 +2.91 100 5 1145.1
Table 9: Step overhead and decay impact summary.
Model Atomic Composite AtomicAvg CompAvg Total Quadrant
gemini-3-pro-preview 62/62 45/45 71.9 83.5 8216.8 High-High
claude-sonnet-4.5 62/62 45/45 65.9 84.0 7868.5 High-High
kimi-k2-thinking 62/62 45/45 71.0 77.9 7904.9 High-High
gpt-5.1 62/62 45/45 66.9 77.7 7642.3 High-High
deepseek-v3.2 62/62 45/45 47.0 90.9 7003.3 High-High
gemini-2.5-flash 62/62 45/45 36.5 82.6 5980.7 Low-High
gpt-oss-120b 62/62 45/45 40.8 81.8 6213.3 Low-High
grok-4-fast 62/62 45/45 46.0 74.6 6210.5 Low-High
gpt-5.1-codex-mini 62/62 45/45 25.4 73.0 5286.3 Low-High
gpt-5 62/62 45/45 56.7 54.3 5959.3 High-Low
claude-haiku-4.5 62/62 45/45 61.5 49.2 6030.7 High-Low
qwen3-235b-thinking 62/62 45/45 57.7 37.6 5270.6 High-Low
qwen3-max 62/62 45/45 44.8 60.0 5472.9 Low-Low
glm-4.6 62/62 45/45 44.7 25.4 3916.1 Low-Low
qwen3-30b-thinking 62/62 45/45 45.9 41.9 4732.0 Low-Low
mimo-v2-flash 62/62 45/45 44.2 49.1 4949.4 Low-Low
qwen3-coder-flash 62/62 45/45 4.7 0.8 327.5 Low-Low
qwen3-coder 62/62 0/45 45.6 0.0 2825.0 Atomic-Only
qwen3-coder-30b 62/62 0/45 48.6 0.0 3015.0 Atomic-Only
devstral-2512 62/62 0/45 46.9 0.0 2905.0 Atomic-Only
Table 10: Model coverage and quadrant classification. Averages are normalized to the 0–100 scale.
F.3 Step Overhead and Decay Impact Summary
Table 9 reports overhead statistics for selected models and the composite score under decay. The overhead statistic is $\Delta=K_{\text{act}}-K_{\text{opt}}$.
Appendix G Case Studies
This section provides representative composite workflows that stress end to end execution. The goal is to highlight failure points such as parameter propagation, prerequisite handling, and protocol sequencing.
G.1 Case 1: Complex Multi step DeFi Workflow
@@
Optimal steps: 6.  Pass rate: 45%.
@@
G.2 Case 2: Batch Approval Workflow
@@
Optimal steps: 3.  Pass rate: 40%.
This pattern stresses multi transaction sequencing and nonce handling. Common failures include nonce conflicts under rapid submissions, gas estimation errors for back to back approvals, and incomplete approvals that allow only one token to proceed.

### v2

@@ D.2 Composite Step Efficiency Decay
D.3 Validator Tolerance Sensitivity
To assess whether the tolerance thresholds influence evaluation outcomes, we analyzed the deviation distribution across 5 rounds $\times$ 20 models $\times$ 107 tasks (10,700 question instances). Table 8 reports the results.
Outcome Count %
Score >0>0, deviation $0\%$ (exact match) 7,013 65.5%
Score =0=0, execution failure / schema error 1,931 18.0%
Score =0=0, transaction revert 493 4.6%
Score =0=0, other failure 1,191 11.1%
Score >0>0, deviation $0\%$ (float precision) 72 0.7%
Table 8: Actual deviation distribution across all 10,700 evaluated question instances (20 models $\times$ 5 rounds $\times$ 107 tasks). Non-zero deviations are all ${\sim}10^{-11}\%$ (floating-point precision), approximately $10^{7}\times$ smaller than the strictest tolerance.
We varied the tolerance multiplier from $0.5\times$ to $3\times$ and re-scored all instances. Under all multiplier settings, all 20 models produce identical scores and rankings (Spearman’s $\rho=1.000$, $\Delta=0\%$). This confirms that tolerances function as safety margins and do not influence reported scores or rankings. The bimodal nature of model outputs — either exactly correct or catastrophically wrong (transaction revert, schema error) — means no question instances fall in the “close but off by a few percent” zone where tolerance thresholds would matter.
@@ Appendix E Additional Analysis Figures
E.1 Total Score Ranking
Figure 6 shows the total score ranking across all 20 models (avg@5).
Figure 6: Model ranking by total score (avg@5, 20 models).
E.2 Composite Pattern Difficulty
@@
E.3 Step Overhead Distribution
@@
E.4 Atomic Subcategory Difficulty
@@ F.1 Quadrant Summary
Table 9 summarizes the four capability profiles based on AtomicAvg and CompositeAvg medians (50.5 and 67.5 respectively).
@@
High-High 7 strong precision and workflows
High-Low 3 strong precision, weaker workflows
Low-High 3 weaker precision, stronger workflows
Low-Low 3 weaker on both splits
Atomic-Only 4 composite split near zero
@@ F.2 Model Coverage and Quadrant Classification
Table 10 reports coverage, normalized averages (avg@5), total score, and the quadrant label based on AtomicAvg and CompositeAvg medians.
Model Atomic Composite AtomicAvg CompAvg Total Quadrant
Claude-Sonnet-4.5 62/62 45/45 69.0 88.0 8235.8 High-High
Gemini-3-Pro 62/62 45/45 69.4 79.1 7863.0 High-High
GPT-5 62/62 45/45 66.6 81.0 7773.7 High-High
GPT-5.1 62/62 45/45 59.9 80.4 7331.7 High-High
Kimi-K2-Thinking 62/62 45/45 58.1 80.7 7238.3 High-High
Gemini-2.5-Flash 62/62 45/45 52.7 83.7 7033.4 High-High
GPT-OSS-120B 62/62 45/45 55.8 77.1 6925.7 High-High
Qwen3-235B 62/62 45/45 58.2 47.0 5724.5 High-Low
Claude-Haiku-4.5 62/62 45/45 52.0 52.6 5588.4 High-Low
Mimo-V2-Flash 62/62 45/45 51.4 38.3 4912.3 High-Low
DeepSeek-V3.2 62/62 45/45 49.5 80.9 6709.2 Low-High
Grok-4-Fast 62/62 45/45 48.1 71.3 6187.0 Low-High
GPT-5.1-Mini 62/62 45/45 40.5 70.9 5700.0 Low-High
Qwen3-Max 62/62 45/45 44.8 64.1 5658.5 Low-Low
GLM-4.6 62/62 45/45 46.5 43.8 4852.0 Low-Low
Qwen3-30B 62/62 45/45 25.1 44.7 3565.5 Low-Low
Devstral-2512 62/62 45/45 46.7 0.5 2913.5 Atomic-Only
Qwen3-Coder 62/62 45/45 29.0 3.6 1961.0 Atomic-Only
Qwen3-Coder-Flash 62/62 45/45 3.6 0.8 260.5 Atomic-Only
Qwen3-Coder-30B 62/62 45/45 2.3 0.1 145.0 Atomic-Only
Table 10: Model coverage and quadrant classification (avg@5). Averages are normalized to the 0–100 scale. Medians: AtomicAvg = 50.5, CompositeAvg = 67.5.
F.3 Step Overhead and Decay Impact Summary
Table 11 reports overhead statistics (avg@5) for models with non-trivial composite scores. The overhead statistic is $\Delta=K_{\text{act}}-K_{\text{opt}}$.
@@ F.3 Step Overhead and Decay Impact Summary
DeepSeek-V3.2 +0.05 44 3 3638.4
Kimi-K2-Thinking +0.17 42 3 3633.3
GPT-5 +0.19 48 3 3645.7
Claude-Sonnet-4.5 +0.34 36 2 3958.0
GPT-OSS-120B +0.41 47 3 3468.1
Gemini-2.5-Flash +0.67 45 3 3768.0
GPT-5.1 +0.70 56 3 3617.1
Gemini-3-Pro +0.83 59 3 3561.0
Grok-4-Fast +1.02 62 3 3206.8
GPT-5.1-Mini +1.24 67 3 3189.8
Qwen3-Max +1.35 63 3 2883.5
Qwen3-Coder +2.36 86 4 160.0
Qwen3-235B +2.47 84 5 2115.9
Claude-Haiku-4.5 +2.51 88 5 2367.4
GLM-4.6 +2.55 89 5 1971.0
Mimo-V2-Flash +2.76 92 5 1723.1
Qwen3-30B +2.88 92 5 2011.5
Table 11: Step overhead and decay impact summary (avg@5). Models with near-zero composite scores (Devstral-2512, Qwen3-Coder-Flash, Qwen3-Coder-30B) are omitted.
Appendix G End-to-End Walkthrough
We illustrate the complete evaluation pipeline with one atomic task: bnb_transfer_percentage.
Step 1: Parameter Instantiation.
The runner selects a natural language template and samples numeric parameters:
  • Template selected: “Transfer {percentage}% of my BNB balance to {recipient}.”
  • Parameters sampled: percentage =15=15, recipient == 0xA1b2...C3d4 (randomly generated).
  • Final instruction: “Transfer 15% of my BNB balance to 0xA1b2…C3d4.”
Step 2: Script Generation.
The model receives the instruction along with system context (chain ID, available libraries, account address, contract ABIs). It generates a TypeScript module that:
  1. 1.
    Queries the current BNB balance via provider.getBalance(account).
  2. 2.
    Computes amount = balance * 15n / 100n.
  3. 3.
    Returns a transaction request: {to: recipient, value: amount}.
Step 3: Execution.
The runner signs and submits the transaction on the forked chain. The transaction executes successfully (receipt status =1=1). The runner records pre-execution and post-execution balances for both the sender and recipient.
Step 4: Validation.
The bnb_transfer_percentage validator performs four checks:
  1. 1.
    Transaction Success (30 pts): receipt status =1=1. ✓
  2. 2.
    Recipient Correct (20 pts): tx.to matches the sampled recipient. ✓
  3. 3.
    Transfer Amount (20 pts): actual transferred amount within 0.1% of expected ($\text{balance}\times 15\%$). ✓
  4. 4.
    Balance Change (30 pts): sender balance decreased by amount ++ gas; recipient balance increased by amount, both within 1% tolerance. ✓
Final score: $30+20+20+30=100$ out of 100.
Key design points.
The expected transfer amount is computed dynamically from the fork state (not hardcoded), so the ground truth is always consistent with the execution environment. The tolerance on balance change (1%) absorbs gas cost estimation differences without admitting incorrect transfers.
Appendix H Case Studies
This section provides representative workflows that illustrate common failure modes. The goal is to highlight failure points such as parameter propagation, prerequisite handling, and protocol sequencing.
H.1 Case 1: Complex Multi step DeFi Workflow
@@
Optimal steps: 6.  Pass rate: 59%.
@@
H.2 Case 2: Batch Approval Workflow
@@
Optimal steps: 3.  Pass rate: 56%.
This pattern stresses multi-transaction sequencing and nonce handling. Common failures include nonce conflicts under rapid submissions, gas estimation errors for back-to-back approvals, and incomplete approvals that allow only one token to proceed.
H.3 Case 3: Natural Language Misinterpretation
Task: bnb_transfer_percentage.
Instruction: “Transfer 15% of my BNB balance to 0xA1b2…C3d4.”
Failure mode: The model interprets “15%” as a fixed amount of 15 BNB rather than 15% of the current balance. This produces a correct transaction that executes successfully but transfers the wrong amount — the validator detects a deviation exceeding the 0.1% tolerance on the Transfer Amount check.
Affected models: Observed in lower-tier models that lack numeric reasoning about percentage-based instructions.
H.4 Case 4: Planning-Level Failure
Task: composite_approve_and_swap.
Instruction: “Swap 0.05 BNB for USDT on PancakeSwap.”
[…]
Impact: The model scores 0 on this task. Step-efficiency decay is irrelevant since the base score is 0.
H.5 Case 5: Partial Execution in Multi-Step Workflows
Task: composite_complete_swap_stake_workflow.
Instruction: “Swap BNB for CAKE, add CAKE-BNB liquidity, then stake the LP tokens.”
Failure mode: The model successfully completes the swap and liquidity addition but fails on the staking step due to an incorrect staking pool address or mismatched LP token. The validator checks the final state (LP tokens staked) and assigns a base score of 0 despite partial progress.
Observation: This pattern accounts for the majority of composite failures among mid-tier models, where individual actions succeed but end-to-end workflow orchestration breaks down.